1. 40 kilometers per hour is equal to how many meters per second?
2. 20 meters per second is equal to many kilometers per hour?

Answer by mayank(15)

1) 40 km/hr = 40*1000/1*60=40000/60=666.67m/sec
2) 20m/sec=20*60/1*1000=1200/1000=1.2km/hr